OpenStudy (naveenbhatia1312):

Convert -630° from degrees to radians. Use the value of π found on a calculator and round answers to four decimal places, as needed. (5 points)

OpenStudy (danjs):

do you remember , a full circle around is 2pi rad

OpenStudy (naveenbhatia1312):

yes

OpenStudy (danjs):

just remember the conversion ratio 360 deg = 2pi rad \[\frac{ -630 ~\deg }{ 1 }*\frac{ 2\pi~rad }{ 360~\deg } = \] multiply correctly to cancel the units

OpenStudy (naveenbhatia1312):

what do you mean

OpenStudy (danjs):

you know the conversion between radians and degrees , it is the same as converting like ft to inches, or anything else...